# HateDay: Insights from a Global Hate Speech Dataset Representative of a Day on Twitter

### v1

@@ 4.3 Feasibility of Hate Speech Moderation
Figure 4: Cost-recall tradeoff in human-in-the-loop moderation. Share of all HateDay tweets flagged as hateful and reviewed by moderators (%) versus share of all hate in HateDay moderated (%). We use the best performing model on HateDay for each language and country (Tables 2 and 3). The dashed line indicates performance of a model that flags tweets as hateful at random.
Most social media platforms prohibit hate speech Singhal et al. (2023) and have done so since their inception Gillespie (2018). In light of our performance results, we investigate the feasibility of hate speech moderation using publicly available hate speech detection models. Specifically, we consider human-in-the-loop moderation, which is an approach used by large social media platforms Avadhanula et al. (2022) whereby content likely to be hateful is flagged by detection models for review by human moderators. We study the trade-off between the amount of human reviewing required and the corresponding share of all hate successfully moderated (Figure 4, details in E.3).
We find that successfully moderating a large share of all hateful content (>80%) using the best publicly available detection models would require human review of least 10% of all tweets posted in a given day. For such a review workload, the total share of hate successfully moderated varies between 70% and 90% for most languages or countries, apart from Turkish where this share is only 40%. 10% of all daily tweets in a given language represents a massive amount of posts, ranging from 19.3 thousand for Kenya to 5.1 million for English. A smaller share of human review would only allow to moderate a small share of all hate: when flagging and reviewing 1% of all tweets posted in a given day, only 20-40% of all hate are moderated for most languages and countries.
@@ 5 Discussion
We find that the proportion of hate speech relative to all social media content is very low, which is in line with past work Gagliardone et al. (2016); Mondal et al. (2017); Park et al. (2022). This contrasts with the high share of social media users who report that they often encounter hate speech in their feeds Vidgen et al. (2019); UNESCO (2023) and points to the potential role of recommendation algorithms in amplifying harmful content Milli et al. (2023).
@@
We find that traditional methods largely overestimate performance in randomly drawn samples of real-world social media data, which is very low, especially for non-European languages. While past work has discussed the risk of such overestimation Arango et al. (2019), due to the biased nature of hate speech datasets Wiegand et al. (2019); Nejadgholi and Kiritchenko (2020), and quantified such risk in the Nigerian Twitter context Tonneau et al. (2024b), we provide evidence of low and overestimated performance across many more linguistic and national contexts. Moreover, our results highlight the need to evaluate NLP tools in realistic settings, especially where human harm might arise. This applies more broadly to the detection of harmful content on social media, such as misinformation Thorne et al. (2018), as well as the evaluation of the bias and safety of generative large language models Ibrahim et al. (2024); Lum et al. (2024); Röttger et al. (2024), both of which have relied so far on non-representative benchmarks for evaluation.
@@
We find that the Perspective API often outperforms other models in hate speech detection across languages, contradicting existing evaluations on traditional hate speech datasets showing that Perspective is outperformed by academic hate speech classifiers Wich et al. (2022); OFCOM (2024). We attribute such dominance to the fact that Perspective is optimized for real-world performance and generalizability to unseen data Lees et al. (2022), contrary to traditional hate speech detection models developed in academia. We also note that the Perspective API is periodically updated and that performance may have improved since the evaluations conducted in past work Pozzobon et al. (2023). Yet, we also observe that Perspective’s dominance does not hold at the country level. This highlights the limits of the language-level approach in developing Perspective and more generally hate speech classifiers Tonneau et al. (2024a), which may not be tailored to national contexts where non-US English and code-mixing is prevalent.
@@
We find that a major reason for the poor performance of hate speech classifiers on HateDay is the high prevalence of offensive content at the top of the hate score distribution, which reduces precision. While offensive content differs from hate speech because it does not target based on the perceived belonging of a particular characteristic, it shares semantic similarities such as the use of swear words and profanities. These similarities likely explain the high hate scores given by classifiers to offensive content. While prior work initially documented this issue Davidson et al. (2017), our findings demonstrate its heightened relevance in social media settings, aggravated by the substantially higher prevalence of offensive content compared to hate speech.
@@
Our findings reveal a misalignment between the targets of hate speech studied in academic work and their real-world prevalence in HateDay. Notably, political-related hate is understudied in academic work compared to its prevalence in HateDay, expanding on past work that points to a low representation of certain targets in hate speech detection resources Yu et al. (2024). We also find that the alignment between academic target coverage and real-world target prevalence has a significant positive correlation with detection performance in HateDay, contrary to the raw amount of detection resources. Improving such alignment, particularly for underrepresented hate types such as political-based hate, is therefore crucial for enhancing real-world detection performance. In the process, political criticism and hate speech must be clearly differentiated to avoid accusations of politically biased moderation Vogels et al. (2020), and hate speech laws being misused for political censorship Strossen (2018).
@@
Our analysis identifies additional challenges for hate speech detection partly identified in past work: First, the difficulty of distinguishing between the use and the mention of hate speech, for instance as part of counter speech Röttger et al. (2022); Gligoric et al. (2024); Jin et al. (2024). Second, the inability of classifiers to capture subtle differences in wording that change a statement from hateful to non-hateful. Third, the lack of context in many posts, especially replies, poses challenges to classification, and calls for more contextualized hate speech detection Pérez et al. (2023).
@@
Our results indicate that hate speech moderation using public hate speech detection models is ineffective at curbing online hate speech at a limited cost. Indeed, low real-world detection performance renders fully automated moderation ineffective and potentially harmful, as it fails to protect users from hate speech while likely censoring non-hateful content, such as counter-speech, confirming concerns from past theoretical work Gorwa et al. (2020). We also show that human-in-the-loop moderation based on public detection models can only effectively moderate a large share of hate speech across our languages and countries of interest at a prohibitively high annotation cost, extending past results for the Nigerian Twitter context Tonneau et al. (2024b) to more languages and countries. Beyond financial costs, the required scale of human review also raises concerns about potential reviewer harm from repeated exposure to harmful content Kirk et al. (2022). Our results also support the claim that detection alone will not solve the hate speech problem Parker and Ruths (2023) and call for complementary solutions to this problem. These include preventive approaches which aim to reshape speech norms and have proven effective to curb hate speech, for instance by prompting users to reconsider harmful posts Katsaros et al. (2022), or by confronting hate speech with counter speech Munger (2017); Hangartner et al. (2021). Finally, we note that our results do not imply that hate speech moderation itself is ineffective, because social media platforms use their own private models for moderation and have access to more data which may improve performance (see Limitations). Yet, without access to real-world performance metrics for platform models, the effectiveness of hate speech moderation on social media remains uncertain.
@@ 6 Conclusion
In this article, we introduced HateDay, the first global hate speech dataset representative of real-world social media settings. We found that evaluations based on previous academic datasets overestimate real-world performance, which is very low. We further discussed the implications of our results for hate speech moderation, namely that it is unfeasible using public detection models. In light of our results, future work is needed to improve hate speech detection performance. In the process, we encourage researchers to ground their evaluation in the real-world settings in which their models may eventually be used. We also encourage platforms to be more transparent on how their detection models perform in social media contexts, in order to better understand the effectiveness of hate speech moderation. We hope that our dataset and experimental results will stimulate work in both of these directions to help solve this important problem.
@@
Low number of positives: The random samples in HateDay used to evaluate hate speech detection in real-world settings contains a low number of hateful examples, ranging from 31 for Kenya to 430 for India. This low number is linked to the very low prevalence of hate speech in our dataset as well as our budget constraint which impeded us from further expanding the annotation effort. While statistically significant, we acknowledge that our performance results on HateDay are necessarily more uncertain, as illustrated by the larger confidence intervals in Tables 5 and 6.
@@
Online hate speech is multimodal: Our work focuses only on text-based hate to limit annotation costs, but we acknowledge that online hate speech is multimodal and that a non-trivial share of this phenomenon, which our analysis necessarily misses, is expressed through images, sounds, or videos Hee et al. (2024).
Limits to representativeness: The tools used to stratify the TwitterDay dataset into language or country-specific sets, namely language detection and user location inference, are imperfect Hecht et al. (2011); Graham et al. (2014); Jurgens et al. (2017). This implies that the representativeness of HateDay may not be perfect, with language or country sets containing some posts in other languages or from other countries. We acknowledge this limitation but argue that HateDay remains as close as a set of representative samples of Twitter as can be with existing stratification tools.
@@
Private data Platforms also rely on several data sources to develop their moderation models, such as the user history, user-graph, conversational context, which we do not have access to. While we consider only public hate speech detection algorithms as a way to flag hate speech in the flow of social media content, we are also aware that platforms use other flagging mechanisms based on private data, such as user reporting or banned word lists, which may increase the share of all hate moderated for a given annotation cost. Finally, the metric platforms aim to reduce is the view count of harmful content such as hate speech, which is arguably more relevant than the sheer prevalence of it from a harm perspective but is also private for the considered timeframe.
@@
The HateDay dataset will be shared publicly upon publication and its intended use is for research purposes only.
@@ Acknowledgements
We thank the annotators for their excellent research assistance. We also thank Dhruti Sanghavi for outstanding visual design work. The study was supported by funding from the United Kingdom’s Foreign Commonwealth and Development Office (FCDO), the World Bank’s Research Support Budget, and the Bill and Melinda Gates Foundation. This work was also supported in part through the NYU IT High Performance Computing resources, services, and staff expertise. The findings, interpretations, and conclusions expressed in this paper are entirely those of the authors. They do not necessarily represent the views of the International Bank for Reconstruction and Development/World Bank and its affiliated organizations or those of the Executive Directors of the World Bank or the governments they represent.
@@ C.1 Language-level Supersets
The language-level supersets are documented in Tonneau et al. (2024a) and obtained from Hugging Face [^3] https://huggingface.co/collections/manueltonneau/hate-speech-supersets-664ef6d2bc40cce7a8b1092f. The number of annotated datapoints in these supersets for each language is:
@@ C.2.1 Surveying approach
To identify HS datasets, we rely on three data sources. First, we inspect the Hate Speech Data Catalogue [^4] https://hatespeechdata.com/ Vidgen and Derczynski (2020). Second, we inspect the datasets listed in the latest survey of HS datasets Poletto et al. (2021). Finally, we conduct a Google search for each country and inspect the links of the first three result pages in each case. We keep only the datasets that fit the following three criteria:
@@
    The dataset focuses on HS, defined broadly as “any kind of communication in speech, writing or behavior, that attacks or uses pejorative or discriminatory language with reference to a person or a group on the basis of who they are, in other words, based on their religion, ethnicity, nationality, race, color, descent, gender or other identity factor” UN (2019).
@@ D.4 Number of parameters
The supervised learning models are largely based on BERT base which has 110 million parameters. The decoder-based models (Aya and Llama3.1) both have 8 billion parameters.
@@ E.2 Individual open-source model performance
egyptian-hate-speech-detection 25.8±0.6 76.9 $\pm$2.0 5.8$\pm$3.0
@@ E.3 Human-in-the-loop moderation
Figure 7: Offensive-to-hateful count ratio versus the share of offensive content in top tweets (in %). Top tweets are defined as the 1% (N=200) top-scored tweets in HateDay using the best performing model on the latter.

### v2

@@ 4.3 Feasibility of Hate Speech Moderation
Figure 4: Cost-recall tradeoff in human-in-the-loop moderation. Share of all HateDay tweets flagged as hateful and reviewed by moderators (%) versus share of all moderated hate in HateDay (%). We use the best model on HateDay for each language and country (Tables 2 and 3). The dashed line indicates performance of a model that flags tweets as hateful at random.
Most social media platforms prohibit hate speech Singhal et al. (2023) and have done so since their inception Gillespie (2018). In light of our poor performance results (§4.1), we investigate the feasibility of hate speech moderation using publicly available hate speech detection models. Given the demonstrated low performance of hate speech detection in real-world social media settings, automatic moderation, whereby content flagged as hateful by a detection model is directly moderated, may be too error-prone. We therefore consider human-in-the-loop moderation, which is a more realistic approach used by large social media platforms Avadhanula et al. (2022) whereby content likely to be hateful is flagged by detection models for review by human moderators. We study the trade-off between the amount of human reviewing required and the corresponding share of all hate successfully moderated (Figure 4, details in §E.3).
We find that successfully moderating a large share of all hateful content (>80%) using the best publicly available detection models would require human review of at least 10% of all daily tweets. For such a review workload, the total share of hate successfully moderated varies between 70% and 90% for most languages or countries, apart from Turkish where this share is only 40%. 10% of all daily tweets may represent a massive amount of posts depending on the context and ranges from 19,300 tweets for Kenya to 5.1 million for English.
A smaller share of human review would lead to most hate being left unmoderated: when flagging and reviewing 1% of all tweets posted in a given day, only 20-40% of all hate is moderated for most languages and countries.
@@ 5 Discussion
We find that the proportion of hate speech relative to all social media content is very low, which is in line with past work Gagliardone et al. (2016); Mondal et al. (2017); Park et al. (2022). This contrasts with recent survey results where two thirds of social media users report that they often encounter hate speech in their feeds UNESCO (2023) and points to the potential role of recommendation algorithms in amplifying harmful content Milli et al. (2025).
@@
We find that traditional evaluation of detection models, on academic datasets and functional tests, largely overestimates performance in randomly drawn samples of real-world social media data. Performance on real-world data is very low, especially for non-European languages. While past work has discussed the risk of such overestimation Arango et al. (2019), due to the biased nature of hate speech datasets Wiegand et al. (2019); Nejadgholi and Kiritchenko (2020), and quantified such risk in the Nigerian Twitter context Tonneau et al. (2024b), we provide evidence of low and overestimated performance across many more linguistic and national contexts. Moreover, our results highlight the need to evaluate NLP tools in realistic settings, especially where human harm might arise. This applies more broadly to the detection of harmful content on social media, such as misinformation Thorne et al. (2018); Magomere et al. (2025), as well as the evaluation of the bias and safety of generative large language models Ibrahim et al. (2024); Lum et al. (2024); Röttger et al. (2024, 2025), both of which have relied so far on non-representative benchmarks for evaluation.
Supervised beats zero-shot
Despite growing enthusiasm for decoder-based large language models (LLMs), we find that they underperform supervised learning for hate speech detection—echoing findings from prior work Nozza (2021). This also reinforces the conclusions of a recent survey of NLP practitioners, which highlights the continued importance of annotated data for maximizing model performance Romberg et al. (2025).
@@
We find that the Perspective API often outperforms other models in hate speech detection across languages, contradicting existing evaluations on traditional hate speech datasets, which showed that Perspective is outperformed by academic hate speech classifiers Wich et al. (2022); OFCOM (2024). We attribute such dominance to the fact that Perspective is optimized for real-world performance and generalizability to unseen data Lees et al. (2022), contrary to traditional hate speech detection models developed in academia. We also note that the Perspective API is periodically updated and that performance may have improved since the evaluations conducted in past work Pozzobon et al. (2023). Still, we find that overall performance remains low, which echoes past evaluations of Perspective’s performance and biases Nogara et al. (2023); Hartmann et al. (2025). Also, we observe that Perspective’s dominance does not always hold at the country level, with the open-source hate speech classifier outperforming Perspective in the case of Nigeria. This may be due to conceptual alignment: the HateDay annotation guidelines were partly based on the same definition and instructions used to train the Nigerian classifier Tonneau et al. (2024b). This also highlights the limits of the language-level approach in developing Perspective and more generally hate speech classifiers Tonneau et al. (2024a), which may not be tailored to national contexts where non-US English and code-mixing is prevalent, as already demonstrated in past work Ghosh et al. (2021); Haber et al. (2023).
@@
We find that a major reason for the poor performance of hate speech classifiers on HateDay is the high prevalence of offensive content at the top of the hate score distribution, which reduces precision. While offensive content differs from hate speech because it does not target an individual or a group based on the perceived belonging of a particular characteristic (e.g., gender, race), it shares lexical features such as the use of swear words and profanities. These similarities likely explain the high hate scores given by classifiers to offensive content. While prior work initially documented this issue Davidson et al. (2017), our findings demonstrate its heightened relevance in real-world social media settings, as it is aggravated by the substantially higher prevalence of offensive content compared to hate speech.
@@
Our findings reveal a misalignment between the targets of hate speech studied in academic work and their real-world prevalence in HateDay. Notably, political-related hate is understudied in academic work compared to its prevalence in HateDay, expanding past work that points to a low representation of certain targets in hate speech detection resources Yu et al. (2024). We also find that the alignment between academic target coverage and real-world target prevalence has a significant positive correlation with detection performance in HateDay, contrary to the raw amount of detection resources. Improving such alignment, particularly for underrepresented hate types such as political-based hate, is therefore crucial for enhancing real-world detection performance. In the process, political criticism and hate speech must be clearly differentiated to avoid accusations of politically biased moderation Vogels et al. (2020) and hate speech laws being misused for political censorship Strossen (2018).
Cross-geographic disparities
Consistent with prior research Röttger et al. (2022a), we find that detection performance is generally lower for non-European languages, such as Arabic, Turkish and Indonesian. Our analysis extends this finding by offering one explanation mentioned in the past paragraph: performance disparities may stem less from varying amounts of annotated training data and more from a misalignment between the target categories prevalent in academic datasets and those that are prevalent in real-world content. This insight may also explain counterintuitive results such as the surprisingly low real-world performance for the English and US contexts and higher performance for India, despite English and the US having the most annotated resources Tonneau et al. (2024a). In real-world English-speaking and US contexts, politically motivated hate is among the most prevalent forms of hate speech; yet, it is underrepresented in US-dominated English academic datasets. By contrast, in the Indian context, there is greater alignment between the types of hate emphasized in academic datasets and those commonly encountered in practice, which may help explain the observed performance differences. Furthermore, the overall prevalence of hate speech is significantly lower in English and US data (0.4% and 0.5% of posts, respectively) compared to Indian data (1.7%), suggesting that our estimate of model performance in English and US settings may be less robust.
@@
We present further challenges for hate speech detection that have partly been identified in past work: First, the difficulty of distinguishing between the use and the mention of hate speech, for instance when slurs or hateful terms are quoted as part of counter speech or educational content Röttger et al. (2021); Gligoric et al. (2024); Jin et al. (2024). Second, the inability of classifiers to capture subtle differences in wording and phrasing that can make a non-hateful statement become hateful Sap et al. (2019). Third, the lack of context in many posts—especially replies—poses challenges to classification and calls for more contextualized hate speech detection Pérez et al. (2023). Our analysis underlines the relevance of these challenges, given their prevalence in real-world settings, and highlights the need to address them as a priority.
@@
Our results indicate that fully automated hate speech moderation using public hate speech detection models is undesirable. Indeed, low real-world detection performance renders fully automated moderation ineffective and potentially harmful, as it fails to protect users from hate speech while likely removing non-hateful content, such as counter-speech. This reinforces concerns raised in earlier theoretical work Gorwa et al. (2020) and complements past empirical findings highlighting the limitations of hate speech detection research for content moderation Ye et al. (2023); Zheng et al. (2024). We also show that human-in-the-loop moderation based on public detection models can effectively moderate a large share of hate speech across our languages and countries of interest but that this implies having humans review a non-trivial share of all daily content, extending past results for the Nigerian Twitter context Tonneau et al. (2024b) to more languages and countries. Beyond financial costs, the required scale of human review also raises concerns about potential reviewer harm from repeated exposure to harmful content Roberts (2016); Kirk et al. (2022). Our results support the claim that detection alone will not solve the hate speech problem Parker and Ruths (2023) and call for complementary solutions. These include preventive approaches that aim to reshape speech norms and have proven effective in curbing hate speech, for instance by prompting users to reconsider harmful posts Katsaros et al. (2022) or by confronting hate speech with counter speech Munger (2017); Hangartner et al. (2021).
@@ 6 Conclusion
In this article, we introduced HateDay, the first global hate speech dataset representative of real-world social media settings. Using HateDay, we showed that evaluating hate speech detection models on standard academic datasets substantially overestimates their real-world performance, which is very low. We explored the implications of this finding for content moderation, concluding that relying on public detection models for automated moderation is currently ill-advised due to their high error rate. Accordingly, improving real-world model performance should be a key focus of future research. We also found that human-in-the-loop moderation can be more accurate, but only if a substantial portion of daily content is manually reviewed—raising important questions about the feasibility and desirability of such an approach at scale, which merits further investigation. Ultimately, we urge researchers to evaluate future detection models within the real-world contexts where they are likely to be deployed. We also call on platforms to provide greater transparency about how their detection systems perform in real social media environments, to better assess moderation effectiveness. We hope that our dataset and findings will drive progress in both areas and contribute to addressing this pressing challenge.
@@
Low number of positives: The random samples in HateDay used to evaluate hate speech detection in real-world settings contains a low number of hateful examples, ranging from 31 for Kenya to 430 for India. This low number is linked to the very low prevalence of hate speech in our dataset as well as our budget constraint, which impeded us from further expanding the annotation effort. While statistically significant, we acknowledge that our performance results on HateDay are necessarily more uncertain, as illustrated by the larger confidence intervals in Tables 5 and 6.
@@
Online hate speech is multimodal: Our work focuses only on text-based hate to limit annotation costs, but we acknowledge that online hate speech is multimodal and that a non-trivial share of this phenomenon, which our analysis necessarily misses, is expressed through images, sounds, or videos Botelho et al. (2021); Hee et al. (2024).
Limits to representativeness: The tools used to stratify the TwitterDay dataset into language or country-specific sets, namely language detection and user location inference, are imperfect Hecht et al. (2011); Graham et al. (2014); Jurgens et al. (2017). This implies that the representativeness of HateDay may not be perfect, with language or country sets containing some posts in other languages or from other countries. We acknowledge this limitation, but argue that HateDay remains the most representative samples of Twitter possible with current stratification tools.
@@
Experiments
Other prompts could lead to different results: We craft a prompt using the terms “hateful” or “non-hateful” (see D.2 for details), which exhibit good performance in past research for hate speech detection using zero-shot learning Plaza-del arco et al. (2023). We do not test other prompts and acknowledge that using other prompts may have an impact on classification performance.
@@
Private data Platforms also rely on several data sources to develop their moderation models, such as the user history, user-graph, conversational context, which we do not have access to. While we consider only public hate speech detection algorithms as a way to flag hate speech in the flow of social media content, we are also aware that platforms use other flagging mechanisms based on private data, such as user reporting or banned word lists, which may increase the share of all hate moderated for a given annotation cost. Finally, the metric platforms aim to reduce is the view count of harmful content such as hate speech, which is arguably more relevant than the sheer prevalence of it from a harm perspective. Unfortunately, such view data is unavailable to us for the considered timeframe.
@@
The intended use of the HateDay dataset is for research purposes only.
@@ Acknowledgements
We thank the annotators for their excellent research assistance. The study was supported by funding from the United Kingdom’s Foreign Commonwealth and Development Office (FCDO), the World Bank’s Research Support Budget, and the Gates Foundation (INV057844). This work was also supported in part through the NYU IT High Performance Computing resources, services, and staff expertise. The findings, interpretations, and conclusions expressed in this article are entirely those of the authors. They do not necessarily represent the views of the International Bank for Reconstruction and Development/World Bank and its affiliated organizations or those of the Executive Directors of the World Bank or the governments they represent. Scott Hale was supported by the ESRC Digital Good Network (grant reference ES/X502352/1). Paul Röttger is a member of the Data and Marketing Insights research unit of the Bocconi Institute for Data Science and Analysis, and is supported by a MUR FARE 2020 initiative under grant agreement Prot. R20YSMBZ8S (INDOMITA).
@@ B.2 Annotation guidelines
We provide a version of the annotation guidelines at https://github.com/manueltonneau/hateday_acl/blob/main/annotation_guidelines.pdf
@@ C.1 Language-level Supersets
The language-level supersets are documented in Tonneau et al. (2024a) and obtained from Hugging Face. [^2] https://huggingface.co/collections/manueltonneau/hate-speech-supersets-664ef6d2bc40cce7a8b1092f The number of annotated datapoints in these supersets for each language is:
@@ C.2.1 Surveying approach
To identify HS datasets, we rely on three data sources. First, we inspect the Hate Speech Data Catalogue [^3] https://hatespeechdata.com/ Vidgen and Derczynski (2020). Second, we inspect the datasets listed in the latest survey of hate speech datasets Poletto et al. (2021). Finally, we conduct a Google search for each country and inspect the links of the first three result pages in each case. We keep only the datasets that fit the following three criteria:
@@
    The dataset focuses on hate speech, defined broadly as “any kind of communication in speech, writing or behavior, that attacks or uses pejorative or discriminatory language with reference to a person or a group on the basis of who they are, in other words, based on their religion, ethnicity, nationality, race, color, descent, gender or other identity factor” UN (2019).
@@ D.4 Number of parameters
The supervised learning models are largely based on the BERT base architecture which has 110 million parameters. The decoder-based models (Aya and Llama3.1) both have 8 billion parameters.
@@ E.2 Individual open-source model performance
egyptian-hate-speech-detection 25.8±0.6 76.9
$\pm$2.0
5.8
$\pm$3.0
@@ E.3 Human-in-the-loop moderation
Figure 7: Offensive-to-hateful count ratio versus the share of offensive content in top tweets (in %). Top tweets are defined as the 1% (N=200) top-scored tweets in HateDay using the best performing model on the dataset.